# Search for gravitationally lensed interstellar transmissions

### v1

@@ II.3 Major noise source
The angular radius of a stellar lens is $R_{\tt L}/z=2.26\times 10^{-9}(R_{\tt L}/R_{\odot})(10\,{\rm pc}/z)\,{\rm rad}$, which is similar to that of the size of the Einstein ring formed by the transmitted beam of light (10). As the angular resolution of the largest telescopes today Turyshev and Toth (2023) is $\lambda/d_{\tt r}\simeq 1\times 10^{-7}(\lambda/1\,\mu{\rm m})(10\,{\rm m}/d_{\tt t})$, they will not be able to resolve neither the star nor the ring, therefore, the use of a coronagraph is out of the question. Thus, the brightness of that lensing star received by the telescope will be the noise that must be dealt with.
@@
Thus, although the stellar flux is very strong (17), it can be significantly reduced by using a narrow bandpass filter at the receiver, centering it at the transmission wavelength (18).
III Detection sensitivity
Quantitative assessment of the signal-to-noise ratio (SNR), is a key to identifying and characterizing received signals. We use results (15) and (18) and estimate the SNR as usual
@@ III Detection sensitivity
Below, we will consider the detection sensitivity in the noise- and signal-dominated regimes.
III.1 Noise-dominated regime
When noise dominates the signal, $Q_{\tt\star}\gg Q_{\tt GL}$, for a receiver close to the optical axis, $\rho\simeq\rho_{\tt GL}$, the maximal value of the SNR is given as
${\rm SNR}_{\tt nd}(\rho)$ $\approx\frac{Q_{\tt ER}}{\sqrt{Q_{\tt\star}}}{}\simeq 287.43\,J^{2}_{0}\big{(}k\theta_{\tt ER}\rho\big{)}\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{\lambda}{1~{}\mu{\rm m}}\Big{)}\Big{(}\frac{10\,{\rm nm}}{\Delta\lambda}\Big{)}^{\frac{1}{2}}\Big{(}\frac{d_{\tt t}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{d_{\tt r}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}\Big{(}\frac{M_{\tt{\tt L}}}{M_{\odot}}\Big{)}\Big{(}\frac{L_{\odot}}{L_{\tt{\tt L}}}\Big{)}^{\frac{1}{2}}\,\sqrt{\frac{t}{1\,{\rm s}}}.$ (20)
Note that (20) was obtained using a modest transmitted power of 1 W. However, that power may be much higher. In fact, modern commercially-available industrial lasers already feature powers of 1–50 kW. Furthermore, coherent beam combination of multiple transmitting laser beams may result in the total transmitted power of up to a few GW and more. In addition, both transmit and receive apertures may be much larger than $\sim$ 1 m which was used in estimates above. In these cases, the signal may be much stronger than the noise, outshining the star, requiring a different evaluation of (19), which will be done next.
III.2 Signal-dominated regime
When signal dominates the noise, $Q_{\tt GL}\gg Q_{\tt L}$, (19) yields
${\rm SNR}_{\tt sd}(\rho)\approx\sqrt{Q_{\tt GL}}$ $\simeq 3.92\times 10^{4}\Big{|}J_{0}\big{(}k\theta_{\tt ER}\rho\big{)}\Big{|}\Big{(}\frac{P_{\tt T}}{1~{}{\rm kW}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}^{\frac{1}{2}}\Big{(}\frac{d_{\tt t}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{d_{\tt r}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\,\sqrt{\frac{t}{1\,{\rm s}}}.~{}~{}~{}~{}~{}$ (21)
Clearly, in a search for transmitted laser signals understanding the SNR is essential. We observe that both estimates (20) and (21) represent handsome levels of detection sensitivity demonstrating feasibility of interstellar power transmission aided by gravitational lensing.
@@ IV Detection strategies
Detection of laser signals encounters several difficulties. The initial challenge is the highly monochromatic nature of lasers emitting light at a specific wavelength. This unique attribute requires finding the exact $\lambda$ matching that of an extraterrestrial signal. This also relates to choosing an appropriate width of the narrow band-pass filter, $\Delta\lambda$, to improve the SNR, especially in the case of the noise-dominated regime (20). There is also directionality of laser transmissions. Contrary to radio waves that can disperse omnidirectionally, laser beams are highly focused and narrow. While interstellar gas and dust are largely permeable to near-IR light interstellar laser signals require precise alignment toward Earth to fall within the detection threshold. As a result: although, stellar brightness (18) and optical background noise are rather well known for each star to be surveyed, the link parameters of the transmitted signal (power, aperture, wavelength, pointing, etc.) used (15) are not available a priory.
Gravitational lensing makes the challenges above less critical, providing a good guidance: The signal that we are looking for has the spatial structure of the PSF of the transmitting lens (9). As such, its structure is known: At typical wavelengths, in the image plane, the spatial frequency of this pattern is on the scale of 100s of meters (as given by (10) and shown in Fig. 3), while its amplitude is proportional to the inverse of the radial distance from the center of the pattern, as seen in (13) as well as in (15)–(16). This is useful for the search.
@@ IV.1 Considering a search campaign
For an interstellar transmission via gravitation lensing, delivery of a transmitter to the focal region of a stellar lens is critical. If that technology is available, then supplying the transmitter with enough propulsion capabilities is similarly plausible. Thus, we may assume that the transmitter is capable of compensating for the relative proper motion between the lens and the Sun, simply directing the signal at the Sun.
@@
The duration over which such lensing event can be observed depends on the width of the received beam and Earth’s orbital velocity. To model the dynamics of the event, we assume a vanishing impact parameter and consider the parallactic motion of an Earth-bound observer, $\mu_{\rm rel}=v_{\oplus}/z$. Under such conditions, the duration of a lensing event, $\Delta t_{\tt ER}$, or the time for the receiver to cross the entire diameter of the Einstein ring (occurring twice a year) is
$\Delta t_{\tt ER}$ $=$ $\frac{2\theta_{\tt ER}}{\mu_{\rm rel}}=\frac{2\sqrt{2r_{g}z_{0}}}{v_{\oplus}}\simeq 14.03\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650~{}{\rm AU}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{30~{}{\rm km/s}}{v_{\oplus}}\Big{)}~{}~{}{\rm hr}.$ (22)
@@
Note that this is the duration of the entire lensing event. A full Einstein ring would only be visible if the impact parameter is very small, and even then, only briefly. In that time, the lens will brighten featuring the amplified laser signal in accord with (9). The actual duration of the appearance of the full Einstein ring depends also on the size of the received the beam, since it corresponds to the observer traversing its width. The spatial structure of the received beam is that of the PSF of the transmitted lens projected on the solar system, with the width of its main lobe given by $\rho_{\tt GL}$ from (11), which is crossed by Earth in just
$\tau_{\tt ER}=\frac{2\rho_{\tt GL}}{v_{\oplus}}\simeq 0.02\,\Big{(}\frac{\lambda}{1~{}\mu{\rm m}}\Big{)}\Big{(}\frac{M_{\odot}}{M_{\tt L}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{650~{}{\rm AU}}{z_{0}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}~{}~{}{\rm s}.$ (23)
@@
The region ±$\pm\rho_{\tt GL}$ where the lensing reaches its maximal value of (9) and detection sensitivities (20) and (21). This is a short but very bright event that can be detected with current technology.
@@
What can be observed a photometric campaign is the varying light amplification as the Earth moves relative to direction to the lensing star with angular separation between them in units of the Einstein ring, $u={\theta}/{\theta_{\tt ER}}$, with $\theta_{\tt ER}$ from (10), which may be expressed as
$u(t,t_{0},u_{0},\Delta t_{\tt ER})=\sqrt{u_{0}^{2}+(t-t_{0})^{2}/\Delta t^{2}_{\tt ER}},$ (24)
@@
where $t_{0}$ is the time of closest alignment and $u_{0}$ is the impact parameter of the event, i.e., the angular separation of the source from the lens at $t_{0}$ expressed in units of $\theta_{\tt ER}$. The quantity $\Delta t_{\tt ER}$ represents the characteristic time scale of the event and is given by (22). The objective is to monitor the brightness of the nearby stars for a signature of a microlensing event. The relevant light amplification factor is characteristic for the lensing within the weak interference region of the lens (Schneider et al., 1992; Gaudi, 2012; Turyshev and Toth, 2023), which is given as below:
@@
$\Delta F(t,t_{0},u_{0},t_{\tt E})=\Big{(}A_{\tt weak.int}(t,t_{0},u_{0},t_{\tt E})-1\Big{)}F_{0},$ (26)
@@ IV.2 Plausible detection strategies
To optimize observations of such transient events, a cooperative telescope network is essential. This network would scan the sky to detect repeatable microlensing events originating at nearby stars with no stellar background that could be responsible. Integrating multiple facilities could enhance spatial resolution, crucial for confident detection and increasing the probability of capturing transmitted signals.
Although photometric measurement capabilities vary among telescopes, any single telescope in the network can perform a broad spectrum of measurements. As a result, a synchronized network of telescopes will be capable of generating detailed imagery of the beam, conducting thorough analyses of its brightness variability.
Although resolving the Einstein rings with the current generation of optical instruments will be challenging, a collaborative network of the astronomical facilities may yield valuable information. With each facility recording brightness measurements and providing its heliocentric position information as a function time, one can use analytical tools available to monitor and study properties of the microlensing event as it unfolds.
@@
We envision using both large telescopes and constellations of small telescopes to capture the photometric data produced by lensing events using the already established photonics and optical engineering technologies. These systems will work in tandem, mutually supplementing observations and offering vital data for predicting events. Several telescopes either in space or both in space and on the ground can be combined in a network to observe the evolving morphology of the Einstein rings as lensing events unfold Turyshev and Toth (2023).
@@
As a result, coordinating ground- and space-based telescope observations is important to detect the structure of the PSF of the transmitting lens as the receiver traverses the projected light field. To do that, evaluating the capabilities of current and upcoming astronomical facilities (see Turyshev and Toth (2023)) including the James Webb Space Telescope, the Nancy Grace Roman Space Telescope, Euclid, the Vera C. Rubin Observatory’s LSST, the Thirty Meter Telescope, and the Extremely Large Telescope, complemented by a network of smaller telescopes is timely, as these facilities may be able to facilitate the search for the faint transient signals transmitted to us by our galactic neighbors.

### v3

@@ III.2 Major noise source
The angular radius of a stellar lens is $R_{\tt L}/z=2.26\times 10^{-9}(R_{\tt L}/R_{\odot})(10\,{\rm pc}/z)\,{\rm rad}$, which is similar to that of the size of the Einstein ring formed by the transmitted beam of light (20). As the angular resolution of the largest telescopes today (see Turyshev and Toth (2023)) is $\lambda/d_{\tt R}\simeq 1\times 10^{-7}(\lambda/1\,\mu{\rm m})(10\,{\rm m}/d_{\tt T})$, they will not be able to resolve neither the star nor the ring, rendering the use of a coronagraph impractical. Consequently, the brightness of the lensing star, as captured by the telescope, constitutes the primary source of noise that needs to be mitigated. Other potential sources of noise, such as the stellar background, the star’s atmosphere, and the interstellar medium, could be considered for a more detailed analysis in future stages, if necessary. However, they are not expected to provide significant contributions.
@@
Thus, despite the strength of the stellar flux (33), its impact can be significantly mitigated by using a narrow bandpass filter at the receiver, as shown in (34). Note that there is already available technology that can be used to search for CW laser signals —specifically, spectrographs with a resolution $R=\lambda/\Delta\lambda\sim 10^{5}$ within the 1–5 $\mu$m bandwidth Kaeufl et al. (2004), which suggests a $\Delta\lambda=0.05$ nm. Clearly, utilizing this specific $\Delta\lambda$ value can further reduce the stellar background noise contribution (34) and enhance detection sensitivity, as detailed by Howard et al. (2004). However, the application of such a narrow bandwidth complicates the search effort for yet unknown signal by presuming prior knowledge of the transmitting wavelength $\lambda$. Consequently, below we will use a broader bandpass filter value used in (34).
III.3 Signal to noise ratio
Quantitative assessment of the SNR, is a key to identifying and characterizing received signals. We use results (32) and (34) and estimate the SNR as usual
@@ III.3 Signal to noise ratio
Below, we will consider the detection sensitivity in the noise- and signal-dominated regimes. As our objective here is to evaluate the feasibility of interstellar power transmissions facilitated by gravitational lensing, to this end, we will rely on straightforward estimates based on the relevant physical parameters, omitting the engineering aspects of the hardware necessary for establishing the laser link (for relevant discussions see Horwath (1996); Howard et al. (2004); Clark and Cahoy (2018).)
When noise dominates the signal, $Q_{\tt\star}\gg Q_{\tt GL}$, the SNR is given as
${\rm SNR}_{\tt nd}(\rho)$ $=\frac{Q_{\tt ER}}{\sqrt{Q_{\tt\star}}}{}\simeq 7.69\,\mu(\rho_{0})\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{L_{\odot}}{L_{\tt{\tt L}}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{10\,{\rm nm}}{\Delta\lambda}\Big{)}^{\frac{1}{2}}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{d_{\tt R}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}\,\sqrt{\frac{t}{1\,{\rm s}}},$ (36)
where we used conservative assumptions on the set of parameters characterizing the interstellar transmission link. Below, we will explore the relevant parameter space that may be lead to improvements in the SNR estimate (36).
We note that the estimate (36) was derived assuming a modest transmission power of 1 W. However, this power could potentially be much higher in practical applications. For instance, modern commercially available industrial lasers already feature powers of 1–50 kW. Additionally, the technique of coherent beam combination—merging multiple transmitting laser beams—could elevate the total transmitted power to several gigawatts (GW) or more. Furthermore, both the transmitting and receiving apertures could be significantly larger than those previously considered. Consequently, the strength of the received signal could vastly surpass that of stellar noise, to the extent of completely overshadowing the star. This scenario necessitates a different approximation method when evaluating (35).
Therefore, in the case, when the signal dominates the noise, $Q_{\tt GL}\gg Q_{\tt L}$, (35) yields
${\rm SNR}_{\tt sd}(\rho)=\sqrt{Q_{\tt GL}}$ $\simeq$ $6.33\times 10^{3}\sqrt{\mu(\rho_{0})}\Big{(}\frac{P_{\tt T}}{1~{}{\rm kW}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{d_{\tt R}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{4}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{4}}\,\sqrt{\frac{t}{1\,{\rm s}}},~{}~{}~{}~{}$ (37)
which is, again, quite conservative, but useful for our purposes. Similarly to (36), result (37) shows the SNR on the optical axis with factor $\mu(\rho_{0})$ modulating the signal for any potential misalignments.
We note that although the estimates (36) and (37) were developed under modest assumptions on the parameters of a transmission link, they indicate high levels of detection sensitivity. These results support the feasibility of projecting significant power across interstellar distances with the assistance of gravitational lensing.
@@ IV Detection strategies
Detecting laser signals presents several challenges. The first hurdle is the highly monochromatic nature of laser light, which emits at a specific wavelength. This characteristic necessitates pinpointing the exact wavelength, $\lambda$, to match an extraterrestrial signal and selecting a suitable narrow band-pass filter width, $\Delta\lambda$, to enhance the SNR, particularly in noise-dominated environments (36). Another challenge is the directionality of laser transmissions. Unlike radio waves, which can spread in all directions, laser beams are highly focused and narrow. Thus, for interstellar laser signals to be detected, they must be precisely aimed towards Earth, especially since interstellar gas and dust, though largely transparent to near-IR light, do not alleviate the need for exact alignment. Consequently, while the stellar brightness (34) and optical background noise are relatively well understood for each nearby star, the specific link parameters of the transmitted signal (32)—such as transmit power, aperture, wavelength, and pointing—are not known in advance.
Gravitational lensing makes some challenges above to be less critical, but some will require consideration: Although the beam of light we are searching for is significantly brighter compared to that of free-space laser transmissions, the transmitted beam arrives in the solar system being highly compressed. Clearly, this feature must be accounted for and appropriately addressed at the transmitter, thus to enhance the probability of signal detection.
@@ IV.1 Considering a search campaign
For interstellar transmissions utilizing gravitational lensing, positioning a transmitter within the focal region of a stellar lens is essential. Assuming the requisite technology for such a deployment exists, it is reasonable to also assume the transmitter would have adequate propulsion capabilities. Consequently, we can infer that the transmitter is designed to counteract the relative proper motion between the lens and the Sun, accurately directing its signal towards the Sun while compensating for the necessary stellar dynamics. This approach, as discussed in Turyshev and Toth (2022b) within the context of exoplanet imaging via the solar gravitational lens (SGL), underscores the importance of integrating advanced propulsion and precise navigational strategies in the transmitter’s design.
As discussed in Sec. II.1, the caustic formed by the annular beam’s propagation around the lens is large. According to (9), its diameter can extend across several AUs, namely $d_{z}\simeq 8.25({\lambda}/{1\,\mu{\rm m}})({1\,{\rm m}}/{d_{\tt T}})({z}/{10\,{\rm pc}})\,{\rm AU}$. Moreover, within the strong interference region, where the intensity of received light peaks, the distance from the optical axis may be as large as $r_{\tt GL}\leq r_{g}(z/z_{0})\simeq 9.37\times 10^{6}(M_{\tt L}/M_{\odot})(z/10\,{\rm pc})(650\,{\rm AU}/z_{0})\,{\rm m}\simeq 1.51R_{\oplus}$. The most intense light, however, is observed when the transmitting telescope aligns directly with the receiver, yielding a received beam diameter notably smaller than in other regions, as discussed in Sec. II.3 and given by (24): $d_{\tt beam}\simeq 3.17\,({z}/{10\,{\rm pc}})({650\,{\rm AU}}/{z_{0}})({d_{\tt T}}/{1~{}{\rm m}})~{}{\rm km}$. These are the three scales where we need to search for the signal.
Therefore, we find that in optical communications utilizing gravitational lenses, precise aiming of the signal transmissions is also crucial. There could be multiple strategies for initiating transmission. For instance, in one scenario, the transmission could be so precisely directed that Earth passes through the targeted spot. Consequently, it’s reasonable to assume that the transmitter would have the capability to track Earth’s movement. Given this precision, one might question whether a deliberately wider beam, capable of encompassing the entire Earth, would be employed instead. This is just few of many scenarios that merit thorough exploration.
As an example, we consider a scenario in which a transmitter targets a specific spot 1 AU from the Sun, allowing Earth to intersect this spot annually. The observability of this lensing event depends on the diameter of the received beam and Earth’s orbital velocity. In a signal-dominated scenario, the SNR, as shown in (37), is influenced by the factor $\sqrt{\mu(\rho_{0})}$, which outside the projected beam diminishes as $({\textstyle\frac{1}{2}}d_{\tt beam}/\rho_{0})^{\frac{1}{2}}$, decreasing slowly with distance from the optical axis. Consequently, even covering the entire Earth, with $\rho_{0}\simeq R_{\oplus}$, the SNR remains exceptionally high at $\sim 6.33\times 10^{3}({\textstyle\frac{1}{2}}d_{\tt beam}/R_{\oplus})^{\frac{1}{2}}\simeq 99.79$, ensuring detectability. Moreover, even at lunar distances, the SNR, estimated at around $12.76$, is sufficiently strong for detection, indicating a robust signal across significant distances.
@@
To model the dynamics of the event, we assume a vanishing impact parameter. Under such conditions, the duration of a lensing event, $\Delta t_{\tt beam}$, or the time for the receiver to cross the entire diameter of the strong interference region is
$\Delta t_{\tt beam}$ $=$ $\frac{2r_{\tt GL}}{v_{\oplus}}\simeq 1.06\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}\Big{(}\frac{z}{10~{}{\rm pc}}\Big{)}\Big{(}\frac{650~{}{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{30~{}{\rm km/s}}{v_{\oplus}}\Big{)}~{}~{}{\rm min}.$ (38)
Depending on the SNR required for signal detection (37), this estimate may be improved. As the radius of the search region increases as $\rho^{\tt SNR}_{0}\sim(6.33\times 10^{3}/{\rm SNR})^{2}{\textstyle\frac{1}{2}}d_{\tt beam}\simeq 6.35\times 10^{8}\,(10/{\rm SNR})^{2}\,{\rm m}$, for ${\rm SNR}=10$ it reaches $\simeq 99.57\,R_{\oplus}$, or nearly 1.7 times of the lunar distance. In this case, the estimate (38) increases to $\Delta t_{\tt beam}\simeq 5.88$ hr. Clearly, as we move away from the optical axis while assuming $P_{\tt T}=1$ kW as in (37), the photon flux (32) drops as $\sim 4.01\times 10^{7}({\textstyle\frac{1}{2}}d_{\tt beam}/\rho_{0})\,{\rm phot/s}$, which at $\rho_{0}=99.57\,R_{\oplus}$, reaches the value of $\sim 1.1\times 10^{2}\,{\rm phot/s}$, all received within the narrow bandwidth of $\Delta\lambda=10$ nm as in (34). This is detectable with the current generation of optical instruments.
We note that increasing the receiver’s aperture to $d_{\tt R}=10$ m in the example above not only increases the flux by a factor of 100 but also improves the SNR by a factor of 10, thereby significantly enhancing the chances of detection. In fact, there is a rather significant parameter space that may be used to develop a number of practical search strategies. Thus, given that the actual transmitter link parameters could exceed those assumed in (37), the peak SNR could be significantly higher. Key variables such as the transmit power ($P_{\tt T}$) and the sizes of the transmit and receive apertures ($d_{\tt T}$ and $d_{\tt R}$, respectively) are prime candidates for upward adjustment. Based on the required minimum SNR, these assumptions will further increase the estimate (38), helping to develop meaningful search strategy.
@@
Note that the maximum brightness would only be observed if the impact parameter is very small, and even then, only briefly. In that time, the lens will brighten featuring the amplified laser signal in accord with (32), assuming signal-dominated transmission. The spatial structure of the received beam is the beam width of $d_{\tt beam}$ from (24), which is crossed by Earth in just
$\tau_{\tt beam}=\frac{d_{\tt beam}}{v_{\oplus}}\simeq 0.11\,\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{30~{}{\rm km/s}}{v_{\oplus}}\Big{)}~{}~{}{\rm s}.$ (39)
@@
The region $d_{\tt beam}$ where the lensing reaches its maximal value of (32) and detection sensitivity (37). This is a short but very bright event that can be detected with current technology, even if receiver is not exactly on the optical axis.
@@
What can be observed a photometric campaign is the varying light amplification as the Earth moves relative to direction to the lensing star with angular separation between them in units of the projected beam spot, $u={\theta}/{\theta_{\tt beam}}$, with $\theta_{\tt beam}=d_{\tt beam}/z$, which may be expressed as
$u(t,t_{0},u_{0},\Delta t_{\tt beam})=\sqrt{u_{0}^{2}+(t-t_{0})^{2}/\Delta t^{2}_{\tt beam}},$ (40)
@@
where $t_{0}$ is the time of closest alignment and $u_{0}$ is the impact parameter of the event, i.e., the angular separation of the source from the lens at $t_{0}$ expressed in units of $\theta_{\tt beam}$. The quantity $\Delta t_{\tt beam}$ represents the characteristic time scale of the event and is given by (38). The objective is to monitor the brightness of the nearby stars for a signature of a microlensing event. The relevant light amplification factor is characteristic for the lensing within the weak interference region of the lens (Schneider et al., 1992; Gaudi, 2012; Turyshev and Toth, 2023), which is given as below:
@@
$\Delta F(t,t_{0},u_{0},\Delta t_{\tt beam})=\Big{(}A_{\tt weak.int}(t,t_{0},u_{0},\Delta t_{\tt beam})-1\Big{)}F_{0},$ (42)
@@ IV.2 Plausible detection strategies
To optimize observations of such transient events, a cooperative telescope network is essential. This network would scan the sky to detect repeatable microlensing events originating at nearby stars with no stellar background that could be responsible. Integrating multiple facilities could enhance spatial resolution, crucial for confident detection and increasing the probability of capturing transmitted signals. Any single telescope in the network can perform a broad spectrum of measurements. As a result, a synchronized network of telescopes will be capable of generating detailed imagery of the beam, conducting thorough analyses of its brightness variability.
@@
Although resolving the Einstein rings with the current generation of optical instruments will be challenging, a collaborative network of astronomical facilities may yield valuable information. With each facility recording brightness measurements and providing its heliocentric position information as a function time, one can use analytical tools available to monitor and study properties of the microlensing event as it unfolds. We envision using both large telescopes and constellations of small telescopes to capture the photometric data produced by lensing events using the already established photonics and optical engineering technologies. These systems will work in tandem, mutually supplementing observations and offering vital data for predicting events. Several telescopes either in space or both in space and on the ground can be combined in a network to monitor the star’s intensity as lensing events unfold Turyshev and Toth (2023).
Generally, the search strategy can be formulated using expression (29). By converting this into a formula for photon flux, expressed as $Q_{\tt R}=(\lambda/hc)P_{\tt R}$, we obtain the following result:
$Q_{\tt R}(\rho_{0})$ $=$ $\Big{(}\frac{\lambda}{hc}\Big{)}P_{\tt T}\frac{\pi({\textstyle\frac{1}{2}}d_{\tt R})^{2}}{\pi(z_{0}+z)^{2}}\Big{(}\frac{d_{\tt T}}{\lambda}\Big{)}^{2}\,\frac{4\sqrt{2r_{g}\tilde{z}}}{d_{\tt T}}\,\mu(\rho_{0}).$ (43)
It is convenient to separate the terms in this expression that relate to the transmitter, the lens and the receiver. For that, for $z\gg z_{0}$, we approximate the impact parameter as $b=\sqrt{2r_{g}\tilde{z}}\simeq R_{\tt L}$, equivalently transforming (43) as below
$Q_{\tt R}(\rho_{0})$ $\simeq$ $\Big{[}\frac{P_{\tt T}d_{\tt T}}{hc}\Big{]}_{\rm transmitter}\Big{[}\frac{R_{\tt L}}{z^{2}}\Big{]}_{\rm lens}\Big{[}\frac{d_{\tt R}^{2}}{\lambda}\Big{]}_{\rm receiver}\mu(\rho_{0}).$ (44)
Clearly, the lens parameters $R_{\tt L}$ and $z$ are known for each target star. Assuming a specific telescope, $d_{\tt R}$ is fixed. With this information, one can conduct a search for a signal $Q_{\tt R}=Q_{\tt R}(\lambda,P_{\tt T}d_{\tt T})$ for a particular transmission wavelength $\lambda$ by varying the combination of the transmitter parameters $P_{\tt T}d_{\tt T}$. Consequently, for each star and a given astronomical facility, for each specific laser wavelength, the search space can be conceptually represented as $Q_{\tt R}=Q_{\tt R}\big{(}\lambda;P_{\tt T}d_{\tt T}\big{)}$. This approach could be used to develop a robust search strategy relying on the capabilities of current technology.
We note that directing an interstellar signal at a specific point in Earth’s heliocentric orbit could lead to false negatives. In such instances, a persistent signal, correlating with Earth’s orbital period, might be observed annually. If this signal is sufficiently weak, it could be mistaken for background noise and disregarded. Therefore, it is critical to observe the targets across multiple spectral bands. Detecting periodic anomalous photometric behavior in one or more channels—especially if multiple laser wavelengths are employed but not uniformly—could signal the presence of the signal of interest. Furthermore, a network of spatially distributed telescopes could facilitate early alerts and enable independent measurements, especially as they will sample the data with different values of factor $\mu(\rho_{0})$.
Once the signal is detected, she spatial distribution of receivers is invaluable, as each will capture a distinct dataset by traveling through the signal along a different path, as encoded $\mu(\rho_{0})$. Correlating the photometric and spectral data from each path enables the reconstruction of the beam’s full profile as it projected onto the solar system. Integrating this information with spectral data from multiple channels reveals the transmitter’s specific features encoded in the beam, such as its power, shape, design, and propulsion capabilities. Additionally, if the optical signal contains encoded information, transmitted via a set of particular patterns, this information will become accessible as well.
@@
As a result, coordinating ground- and space-based telescope observations is important to detect the structure of the PSF of the transmitting lens as the receiver traverses the projected light field. To do that, evaluating the capabilities of current and upcoming astronomical facilities (see Turyshev and Toth (2023) for details) including the James Webb Space Telescope, the Nancy Grace Roman Space Telescope, Euclid, the Vera C. Rubin Observatory’s LSST, the Thirty Meter Telescope, and the Extremely Large Telescope, complemented by a network of smaller telescopes is timely, as these facilities may be able to facilitate the search for the faint transient signals transmitted to us by our galactic neighbors.
Although we considered the case where transmitter points at a spot at 1 AU from the Sun, other transmission scenarios exist, including one that involves directly targeting Earth and compensating for its orbital motion Howard et al. (2004). This particular approach would necessitate enhanced propulsion and targeting capabilities for the transmitting spacecraft, thereby adding complexity to the design of both the spacecraft and its mission. Additionally, in such a scenario, the lensing star would display additional brightness due to persistent lensing effect, further complicating detection efforts if only a single telescope is employed. However, a coordinated network of ground- and space-based telescopes, discussed here, will be capable of effectively manage this transmission scenario thus improving chances of signal detection. Furthermore, monitoring the target star across various spectral channels and from diverse spatial positions enhances the chances of detecting unusual photometric variability that might manifest in one or several narrow spectral bands.